OpenStudy (s117):

what is 2.64 as a fraction?

OpenStudy (s117):

@Directrix

OpenStudy (anonymous):

Expressed as a top-heavy fraction in its simplest form, 2.64 is equal to 66/25 or sixty-six over twenty-five. Expressed as a mixed number in its simplest form, 2.64 is equal to 2 16/25 or two and sixteen twenty-fifths medal ?

OpenStudy (anonymous):

2.64= 2 64/100=(2*100+64)/100=264/100=66/25

Directrix (directrix):